100 ml of 0.6 N H2SO4 and 200 ml of 0.3N HCL were mixed together.what is normality of resulting solution
 

Grade:11

1 Answers

Bolloju anilkumar
5 years ago
N1V1+N2V2=NVtotal
100×0.6+200×0.3=N×300(N1=0.6N,N2=0.3N)
60+60=N×300(Vtotal=300ml)
120=N×300
N=120/300
N=0.4N